OpenStudy (anonymous):

Jon, John and Jan worked together to type a manuscript in 3 hours. If Jon and Jan are to work together, they can finish typing the same manuscript in 4 hours. On the other hand, if john worked alone for 7 hours and then is joined by Jan, they can finish typing it in two more hours. How many hours will each need to type the manuscript alone?

hero (hero):

So basically, you have to create 3 separate equations in order to solve

OpenStudy (anonymous):

I've done that, but I'm not sure how to solve them.

hero (hero):

The third one is a trick

hero (hero):

You probably don't have the same equations as me. What did you use as variables?

hero (hero):

Let me see your equations

hero (hero):

@satellite73

hero (hero):

For some reason, I was getting these three equations using your method:

OpenStudy (anonymous):

i have a headache already

hero (hero):

a = Jon b = John c = Jan \[\large \frac{abc}{ab + bc + ac} = 3\] \[\large \frac{ac}{a + c} = 4\] \[\large \frac{bc}{b+ c} = 2\]

OpenStudy (anonymous):

\[r_1+r_2+r_3=\frac{1}{3}\] \[r_1+r_3=\frac{1}{4}\] \[r_2=\frac{1}{7}\]

hero (hero):

But apparently there's something wrong with my last equation, but I don't know why

OpenStudy (anonymous):

since \(r_2=\frac{1}{7}\) we get the first equation is ... no no that was wrong

hero (hero):

I posted equations of my own satellite. Look at those

OpenStudy (anonymous):

\(r_2\) is not \(\frac{1}{7}\)

hero (hero):

Do you see the ones I posted above?

OpenStudy (anonymous):

yeah, i think the last one is wrong last statement is tricky

OpenStudy (anonymous):

this one is tricky if john worked alone for 7 hours and then is joined by Jan, they can finish typing it in two more hours i screwed up with mine

hero (hero):

Mine is even more messed up

hero (hero):

I have no clue how to write it

OpenStudy (anonymous):

i think your first two are right. last one is wrong

hero (hero):

That's why I brought you here. To help me figure out the last one

OpenStudy (anonymous):

ok hold on, i think we can get second rate

hero (hero):

It uses your methods afterall

OpenStudy (anonymous):

well this one is a bit different because of the mix at the end

OpenStudy (anonymous):

lets see if slow is good first statement Jon, John and Jan worked together to type a manuscript in 3 hours so \(a+b+c=\frac{1}{3}\)

OpenStudy (anonymous):

If Jon and Jan are to work together, they can finish typing the same manuscript in 4 hours tells us \(a+c=\frac{1}{4}\) and so we can solve for \(b\)

OpenStudy (anonymous):

this tells us \(b=\frac{1}{12}\)

OpenStudy (anonymous):

we get John needs 12 hours to do the job by himself, now lets solve for Jan

hero (hero):

How come you're not using the \(\frac{MN}{M + N}\)

OpenStudy (anonymous):

because that works when question says "how long will it take them both" given the rate of each i think your first and second equations are correct, but the third one is wrong and besides, we need the rate of each, so it makes more sense to solve for them individually

OpenStudy (anonymous):

at any rate, (no pun intended) we got John's rate at \(\frac{1}{12}\) and using that we can get Jan's rate

OpenStudy (anonymous):

john works for 7 hours, completes \(\frac{7}{12}\) of the job, leaving \(\frac{5}{12}\) to complete, then it takes two hours with the additional rate of jan, so \[(\frac{1}{12}+c)\times 2=\frac{5}{12}\]

OpenStudy (anonymous):

\[\frac{2}{12}+2c=\frac{5}{12}\] \[2c=\frac{3}{12}\] \[c=\frac{1}{8}\] by this time my arithmetic and algebra are probably suspect

OpenStudy (anonymous):

and finally \[a+\frac{1}{12}+\frac{1}{8}=\frac{1}{3}\] and i guess i should actually solve because now i am curious

hero (hero):

I already knew what the answers were.

hero (hero):

Just couldn't figure out the equation for the last one

OpenStudy (anonymous):

you knew? lol ok did you get 8 12 and 8?

hero (hero):

Yeah

OpenStudy (anonymous):

the problem with that slick method i like so much is that it doesn't work in exactly these types of problems, where one person works for a while and then the other joins that is because the right hand side is no long one (one job) but something like five twelfths of a job

hero (hero):

Oh, I see what you mean

hero (hero):

But when you did it, you only used the first two equations as well.

hero (hero):

So my way should work for that too

OpenStudy (anonymous):

yeah maybe, but how much of the job has one completed before the other joins? you need to know how much is left i bet you can come up with an equation in \(b\) and \(c\) that would give you that, but i don't know it

OpenStudy (anonymous):

it would be rather complicated i think

hero (hero):

If anyone can, @satellite73 can

OpenStudy (anonymous):

thanks for the vote of confidence now it is time for the twilight zone

hero (hero):

I'm disappointed that there's no way to find b using the method I chose